OpenStudy (anonymous):

8x + 7 − 2x = 27 − 2

OpenStudy (anonymous):

\[8x+7-2x=27-2\]Start by combining like terms; \[6x+7=25\]Now subtract 7 from both sides; \[6x=18\]Now divide both sides by 6; \[x=3\]

OpenStudy (anonymous):

Okay, I understand now, thanks :)

OpenStudy (anonymous):

No problem. (: